OpenStudy (anonymous):

For the 1 -5, using the given. matrices find the value of its determinant: |2 3 | |-1 4| Choose one answer. a. 11 b. -22 c. 4 d. 5

OpenStudy (anonymous):

i got 9 but its wrong obviously

OpenStudy (anonymous):

Do you know how to find the determinant?

OpenStudy (anonymous):

yes i think i have examples in front of me

OpenStudy (jackellyn):

\[A=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] det(A)=a(d)-c(b) can you apply this to your matrix \[\left[\begin{matrix}2 & 3 \\ -1 & 4\end{matrix}\right]\]?

OpenStudy (anonymous):

i did tht and i got 9

OpenStudy (anonymous):